# Universal Time-Series Representation Learning: A Survey

### v1

@@ 1. Introduction
1.1. Background
A time series is a sequence of data points recorded in chronological order, reflecting the complex dynamics of particular variables or phenomena over time. Time-series data can represent various meaningful information across application domains at different time points, enabling informed decision-making and predictions, such as sensor readings in the Internet of Things (Fathy et al., 2018; Cook et al., 2019), measurements in cyber-physical systems (Giraldo et al., 2018; Luo et al., 2021), fluctuation in stock markets (Ge et al., 2022; Cao, 2022), and human activity on wearable devices (Chen et al., 2021a; Gu et al., 2021). However, to extract and understand meaningful information from such complicated observations, we need a mechanism to represent these time series, which leads to the emergence of time-series representation research. Based on the new representations, we can effectively perform various downstream time-series analyses (Esling and Agon, 2012), e.g., forecasting (Lim and Zohren, 2021), classification (Ismail Fawaz et al., 2019), regression (Tan et al., 2021a), and anomaly detection (Choi et al., 2021). Fig. 1 depicts the basic concept of representation methods for time-series data.
@@ 1.1. Background
Early attempts (Sharma et al., 2020) represent time series using piecewise linear methods (e.g., piecewise aggregate approximation), symbolic-based methods (e.g., symbolic aggregate approximation), feature-based methods (e.g., shapelets), or transformation-based methods (e.g., discrete wavelet transform). These traditional time-series representation methods are known to be time-consuming and less effective because of their dependence on domain knowledge and the poor generality of predefined priors. Since the quality of representations significantly affects the downstream task performance, many studies propose to learn the meaningful time-series representations automatically (Zerveas et al., 2021; Yue et al., 2022; Yang and Hong, 2022). The main goal of these studies is to obtain high-quality learned representations of time series that capture valuable information within the data and unveil the underlying dynamics of the corresponding systems or phenomena. Among several approaches, neural networks or deep learning (DL) have demonstrated unprecedented performance in extracting hidden patterns and features from a wide range of data, including time series, without the requirement of manual feature engineering.
Given the sequential nature of time series, recurrent neural networks (RNN) and their variants, such as long short-term memory and gated recurrent unit, are considered a popular choice for capturing temporal dependencies in time series (Lalapura et al., 2021; Mao and Sejdić, 2022). Nevertheless, recurrent-based networks are complex and computationally expensive. Another line of work adopts one-dimensional convolutional neural networks (CNN) to improve the computational efficiency with the parallel processing of convolutional operations (Iwana and Uchida, 2021). Even though RNN- and CNN-based models are shown to be good at capturing temporal dependencies, they cannot explicitly model the relationship between different variables within a multivariate time series. Many studies propose to use attention-based networks or graph neural networks to jointly learn the temporal dependencies in each variable and the correlations between different variables in multivariate time series using attention mechanisms or graph structures (Wen et al., 2023; Jin et al., 2023a). Despite the significant progress in architectural design, time series can be collected irregularly or have missing values caused by sensor malfunctions in real-world scenarios, making the commonly used neural networks inefficient due to the adversarial side-effect during the imputation process. Consequently, recent research integrates neural ordinary differential equations into existing networks such that the models can produce continuous hidden states, thereby being robust to irregular time series (Rubanova et al., 2019a; Sun et al., 2020).
@@
In addition, the reliability and efficacy of DL-based methods are generally contingent upon the availability of sufficiently well-annotated data, commonly known as supervised learning. Time-series data, however, is naturally continuous-valued, contains high levels of noise, and has less intuitively discernible visual patterns. In contrast to human-recognizable patterns in images or texts, time-series data can have inconsistent semantic meanings in real-world settings across application domains. As a result, obtaining a well-annotated time series is inevitably inefficient and considerably more challenging even for domain experts due to the convoluted dynamics of the time-evolving observations collected from diverse sensors or wearable devices with different frequencies. For example, we can collect a large set of sensor signals from a smart factory, while only a few of them can be annotated by domain experts. To circumvent the laborious annotation process and reduce the reliance on labeled instances, there has been a growing interest in unsupervised and self-supervised learning approaches using self-generated labels from various pretext tasks without relying on human annotation (Eldele et al., 2023; Zhang et al., 2023a; Ma et al., 2023; Meng et al., 2023).
@@
While unsupervised and self-supervised representation learning share the same objective to extract latent representations from intricate raw time series without the human-annotated labels, their underlying mechanism is different. Unsupervised learning methods (Meng et al., 2023) usually adopt autoencoders and sequence-to-sequence models to learn meaningful representations using reconstruction-based learning objectives. However, accurately reconstructing the complex time-series data remains challenging, especially with high-frequency signals. On the contrary, self-supervised learning methods (Zhang et al., 2023a) leverage pretext tasks to autonomously generate labels by utilizing intrinsic information derived from the unlabeled data. Lately, pretext tasks with contrasting loss (also known as contrastive learning) have been proposed to enhance learning efficiency through discriminative pre-training with self-generated supervised signals. Contrastive learning aims to bring similar samples closer while pushing dissimilar samples apart in the feature space. These pretext tasks are self-generated challenges the model learns to solve from the unlabeled data, thereby being able to produce meaningful representations for multiple downstream tasks (Ericsson et al., 2022).
@@
Figure 2. Key design elements and downstream evaluation protocols.
@@ 1.2. Related Surveys
According to the background discussed above, there are three fundamental elements (also illustrated in Fig. 2) in designing a state-of-the-art universal representation learning method for time series: neural architectures, learning objectives, and training data. The neural architectures are designed to capture underlying temporal dependencies in time series and inter-relationships between variables of multivariate time series. Also, one or multiple learning objectives (i.e., loss functions) are newly defined to learn high-quality representations. These learning objectives are sometimes called pretext tasks if pseudo labels are generated. Last, to enhance the utility and quality of available training data, data-related techniques (e.g., augmentation) are employed or introduced.
Despite having the three key design elements, most existing surveys for time-series representation learning review the literature exclusively on either the neural architectural aspects (Längkvist et al., 2014) or the learning aspects (Zhang et al., 2023a; Ma et al., 2023; Meng et al., 2023). An early related survey article by Längkvist et al. (2014) reviews the deep learning for time-series data as unsupervised feature learning algorithms. The survey focuses particularly on neural architectures with a discussion on classical time-series applications. After about a decade, a few survey papers review time-series representation learning methods by focusing on learning objective aspects. For example, Zhang et al. (2023a) and Deldari et al. (2022) review self-supervised learning-based models, while, for a broader scope, Meng et al. (2023) review unsupervised learning-based methods. Similarly, Ma et al. (2023) present the survey for any learning objectives with a focus on analyzing reviewed articles from transfer learning and pre-training perspectives. For a smaller scope, a survey (Eldele et al., 2023) reviews the state-of-the-art studies that specifically tackle label scarcity in time-series data. Unlike existing survey papers, this paper comprehensively reviews the representation learning methods for time series by focusing on their universality with discussions on their intuitions and insights into how these methods enhance the quality of learned representations from all three design aspects. Specifically, we aim to review and identify research directions on how recent state-of-the-art studies design the neural architecture, devise corresponding learning objectives, and utilize training data to enhance the quality of the learned representations of time series for downstream tasks. Table 1 summarizes the differences between our survey and the related work.
@@
Survey Article Novel Taxonomy Key Design Elements in Review Survey Coverage Future Research Direction
Neural
Architecture
Learning
Objective
Training
Data
Universality Irregularity
Längkvist et al. (2014) $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$
Deldari et al. (2022) $\checkmark$ $\times$ Only Self-Supervised Learning $\times$ $\checkmark$ $\times$ $\checkmark$
Eldele et al. (2023) $\checkmark$ $\times$ $\checkmark$ Only Augmentation $\checkmark$ $\times$ $\checkmark$
Ma et al. (2023) $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$
Zhang et al. (2023a) $\checkmark$ $\times$ Only Self-Supervised Learning $\times$ $\times$ $\times$ $\checkmark$
Meng et al. (2023) $\checkmark$ $\times$ Only Unsupervised Learning $\times$ $\checkmark$ $\times$ $\checkmark$
Our Survey $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ 1.3. Survey Scope and Literature Collection
1.3.1. Keywords
“time series” AND “representation”, “time series” AND “embedding”, “time series” AND “encoding”, “time series” AND “modeling”, “time series” AND “deep learning”, “temporal” AND “representation”, “sequential” AND “representation”, “audio” AND “representation”, “sequence” AND “representation”, and (“video” OR “action”) AND “representation”. We use these keywords to search well-known repositories, including ACM Digital Library, IEEE Xplore, Google Scholars, Semantic Scholars, and DBLP, for the relevant papers.
1.3.2. Inclusion Criteria
The initial set of papers found with the above search queries are further filtered by the following criteria. Only papers meeting the criteria are included for review.
@@ 1.3.2. Inclusion Criteria
1.3.3. Quantitative Summary of Selected Papers
Given the above keywords and inclusion criteria, we selected 105 papers in total. Fig. 3 shows the quantitative summary of the paper selected for review. We can notice from Fig. 2(a) that neural architectures and learning objectives are similarly considered important in designing state-of-the-art methods. Most papers were published at NeurIPS, followed by AAAI, CVPR, and ICLR (Fig. 2(b)). According to Fig. 2(c), we expect more papers on this topic will be published in the future.

### v3

@@ 1. Introduction
A time series is a sequence of data points recorded in chronological order, reflecting the complex dynamics of particular variables or phenomena over time. Time-series data can represent various meaningful information across application domains at different time points, enabling informed decision-making and predictions, such as sensor readings in the Internet of Things (Fathy et al., 2018; Shin et al., 2019; Trirat et al., 2023), measurements in cyber-physical systems (Giraldo et al., 2018; Luo et al., 2021), fluctuation in stock markets (Ge et al., 2022; Cao, 2022), and human activity on wearable devices (Chen et al., 2021c; Gu et al., 2021). However, to extract and understand meaningful information from such complicated observations, we need a mechanism to represent these time series, which leads to the emergence of time-series representation research. Based on the new representations, we can effectively perform various downstream time-series analyses (Esling and Agon, 2012), e.g., forecasting (Lim and Zohren, 2021), classification (Ismail Fawaz et al., 2019), regression (Tan et al., 2021), and anomaly detection (Choi et al., 2021). Fig. 1 depicts the basic concept of representation methods for time-series data.
Early attempts (Sharma et al., 2020) represent time series using piecewise linear methods (e.g., piecewise aggregate approximation), symbolic-based methods (e.g., symbolic aggregate approximation), feature-based methods (e.g., shapelets), or transformation-based methods (e.g., discrete wavelet transform). These traditional time-series representation methods are known to be time-consuming and less effective because of their dependence on domain knowledge and the poor generality of predefined priors. Since the quality of representations significantly affects the downstream task performance, many studies propose to learn the meaningful time-series representations automatically (Deldari et al., 2022). The main goal of these studies is to obtain high-quality learned representations of time series that capture valuable information within the data and unveil the underlying dynamics of the corresponding systems or phenomena. Among several approaches, neural networks or deep learning (DL) have demonstrated unprecedented performance in extracting hidden patterns and features from a wide range of data, including time series, without the requirement of manual feature engineering. Note that representation learning can be categorized into task-specific (Cheng et al., 2023) and task-agnostic(Yue et al., 2022; Liu and Chen, 2023), i.e., universal, approaches. This survey focuses exclusively on universal representation learning, which refers to methods that are designed for and evaluated on at least two downstream tasks.
Given the sequential nature of time series, recurrent neural networks (RNNs) and their variants, such as long short-term memory and gated recurrent unit, are considered a popular choice for capturing temporal dependencies in time series (Lalapura et al., 2021). Nevertheless, RNNs are complex and computationally expensive. Another line of work adopts one-dimensional convolutional neural networks (CNN) to improve the computational efficiency with the parallel processing of convolutional operations (Iwana and Uchida, 2021). Even though RNN- and CNN-based models are shown to be good at capturing temporal dependencies, they cannot explicitly model the relationship between different variables within a multivariate time series. Many studies propose to use attention-based networks or graph neural networks to jointly learn the temporal dependencies in each variable and the correlations between different variables using attention mechanisms or graph structures (Wen et al., 2023; Jin et al., 2023a). Despite the significant progress in architectural design, time series can be collected irregularly or have missing values caused by sensor malfunctions in real-world scenarios, making the commonly used neural networks inefficient due to the adversarial side-effect during the imputation process. Consequently, recent research integrates neural differential equations into existing networks such that the models can produce continuous hidden states, thereby being robust to irregular time series (Rubanova et al., 2019a; Sun et al., 2020).
@@
[…]
Basic concept of time-series representation methods.
@@
The reliability and efficacy of DL-based methods are generally contingent upon the availability of sufficiently well-annotated data, commonly known as supervised learning. Time-series data, however, is naturally continuous-valued, contains high levels of noise, and has less intuitively discernible visual patterns. In contrast to human-recognizable patterns in images or texts, time series can have inconsistent semantic meanings in real-world settings across application domains. As a result, obtaining a well-annotated time series is inevitably inefficient and considerably more challenging even for domain experts due to the convoluted dynamics of the time-evolving observations collected from diverse sensors or wearable devices with different frequencies. For example, we can collect a large set of sensor signals from a smart factory, while only a few of them can be annotated by domain experts. To circumvent the laborious annotation process and reduce the reliance on labeled instances, there has been a growing interest in unsupervised and self-supervised learning using self-generated labels from various pretext tasks without relying on human annotation (Eldele et al., 2023).
@@
While unsupervised and self-supervised representation learning share the same objective to extract latent representations from intricate raw time series without the human-annotated labels, their underlying mechanism is different. Unsupervised learning methods (Meng et al., 2023) usually adopt autoencoders and sequence-to-sequence models to learn meaningful representations using reconstruction-based learning objectives. However, accurately reconstructing the complex time-series data remains challenging, especially with high-frequency signals. On the contrary, self-supervised learning methods (Zhang et al., 2023e) leverage pretext tasks to autonomously generate labels by utilizing intrinsic information derived from the unlabeled data. Lately, pretext tasks with contrasting loss (also known as contrastive learning) have been proposed to enhance learning efficiency through discriminative pre-training with self-generated supervised signals. Contrastive learning aims to bring similar samples closer while pushing dissimilar samples apart in the feature space. These pretext tasks are self-generated challenges the model learns to solve from the unlabeled data, thereby being able to produce meaningful representations for multiple downstream tasks (Ericsson et al., 2022).
@@
[…]
Key design elements and evaluation protocols of a time-series representation learning framework.
Figure 2. Key design elements and evaluation protocols of a time-series representation learning framework.
@@ 1.1. Related Surveys
According to the background discussed above, there are three fundamental elements (also illustrated in Fig. 2) in designing a state-of-the-art universal representation learning method for time series: training data, network architectures, and learning objectives. To enhance the utility and quality of available training data, data-related techniques (e.g., augmentation) are employed or introduced. The neural architectures are then designed to capture underlying temporal dependencies in time series and inter-relationships between variables of multivariate time series. Last, one or multiple learning objectives (i.e., loss functions) are newly defined to learn high-quality representations. These learning objectives are sometimes called pretext tasks if pseudo labels are generated.
Despite having the three key design elements, most existing surveys for time-series representation learning review the literature exclusively on either the neural architectural aspects (Längkvist et al., 2014; Wang et al., 2024a) or the learning aspects (Zhang et al., 2023e; Ma et al., 2023; Meng et al., 2023). An early related survey article by Längkvist et al. (2014) review the DL for time series as unsupervised feature learning algorithms. The survey focuses particularly on neural architectures with a discussion on classical time-series applications. After about a decade, a few survey papers review time-series representation learning methods by focusing on learning objective aspects. For example, Zhang et al. (2023e) and Deldari et al. (2022) review self-supervised learning-based models, while, for a broader scope, Meng et al. (2023) review unsupervised learning-based methods. Similarly, Ma et al. (2023) present the survey for any learning objectives with a focus on analyzing reviewed articles from transfer learning and pre-training perspectives. For a smaller scope, a survey (Eldele et al., 2023) reviews the state-of-the-art studies that specifically tackle label scarcity in time-series data. With the advent of foundation and large language models (LLMs), recent articles (Ye et al., 2024; Liang et al., 2024) review the adaptation of these models to time series with the main focus on learning aspects. Unlike these papers, we comprehensively review the representation learning methods for time series by focusing on their universality—effectiveness across diverse downstream tasks—with discussions on their intuitions and insights into how these methods enhance the quality of learned representations from all three design aspects. Specifically, we aim to review and identify research directions on how recent state-of-the-art studies design the neural architecture, devise corresponding learning objectives, and utilize training data to enhance the quality of the learned representations of time series for downstream tasks. Table 1 summarizes the differences between our survey and the related work.
@@
Survey Article Key Design Elements in Review Survey Coverage
Training Data Neural Architecture Learning Objective Universality Irregularity Experimental Design
Generation Augmentation Supervised Unsupervised Self-Supervised
Längkvist et al. (2014) $\checkmark$
Deldari et al. (2022) $\checkmark$ $\checkmark$
Eldele et al. (2023) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Ma et al. (2023) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Zhang et al. (2023e) $\checkmark$ $\checkmark$
Meng et al. (2023) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Ye et al. (2024) $\checkmark$ Only LLMs $\checkmark$ $\checkmark$ $\checkmark$
Liang et al. (2024) $\checkmark$ $\checkmark$ $\checkmark$
Wang et al. (2024a) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Our Survey $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ 1.2. Survey Scope and Literature Collection
Keywords. “time series” AND “representation”, “time series” AND “embedding”, “time series” AND “encoding”, “time series” AND “modeling”, “time series” AND “deep learning”, “temporal” AND “representation”, “sequential” AND “representation”, “audio” AND “representation”, “sequence” AND “representation”, and (“video” OR “action”) AND “representation”. We use these keywords to search well-known repositories, including ACM Digital Library, IEEE Xplore, Google Scholars, Semantic Scholars, and DBLP, for the relevant papers.
Inclusion Criteria. The initial set of papers found with the above search queries are further filtered by the following criteria. Only papers meeting the criteria are included for review.
@@
Table 2. The proposed taxonomy of universal representation learning for time series.
Design Element Coarse-to-Fine Taxonomy References
Data-Centric Approaches (Training Data) Improving Quality (§$\S$3.1) Sample Selection (Chen et al., 2021a; Franceschi et al., 2019)
Time-Series Decomposition (Zeng et al., 2021; Behrmann et al., 2021; Wang et al., 2018; Yang et al., 2022a; Fang et al., 2023)
Input Space Transformation (Anand and Nayak, 2021; Biloš et al., 2022; Lee et al., 2022a; Wu et al., 2023; Zhong et al., 2023; Xu et al., 2024)
Increasing Quantity (§$\S$3.2) Augmentation Random Augmentation (Yue et al., 2022; Zhang et al., 2022b, 2023f; Liu and Chen, 2023)
Policy-Based Augmentation (Kim et al., 2023c; Chen et al., 2022; Kim et al., 2023b; Zhang and Crandall, 2022; Aboussalah et al., 2023; Luo et al., 2023; Yang et al., 2022b; Yang and Hong, 2022; Shin et al., 2023; Demirel and Holz, 2023; Zheng et al., 2024; Duan et al., 2024; Li et al., 2024)
Generation and Curation (Nguyen et al., 2023; Zhao et al., 2023; Liu et al., 2024; Goswami et al., 2024)
Neural Architectural Approaches (Network Architecture) Task-Adaptive Submodules (§$\S$4.1) (Chen et al., 2021a; Wang et al., 2023c; Gorbett et al., 2023; Liang et al., 2023a; Gao et al., 2024; Zhang et al., 2024; Goswami et al., 2024)
General Temporal Modeling (§$\S$4.2) Intra-Variable Modeling Long-Term Modeling (Zerveas et al., 2021; Tonekaboni et al., 2022; Zhang et al., 2023c; Nguyen et al., 2023; Luo and Wang, 2024; Han et al., 2020; Franceschi et al., 2019)
Multi-Resolution Modeling (Liu et al., 2022; Zhong et al., 2023; Fraikin et al., 2023; Nguyen et al., 2023; Liu et al., 2020; Sanchez et al., 2019; Wang et al., 2023a; Sener et al., 2020; Wang et al., 2024c; Eldele et al., 2024)
Inter-Variable Modeling (Zhong et al., 2023; Xie et al., 2022; Guo et al., 2021; Cai et al., 2021; Xiao et al., 2023; Wang et al., 2023b, 2024c; Luo and Wang, 2024; Zhang et al., 2024; Wang et al., 2024b)
Frequency-Aware Modeling (Yang and Hong, 2022; Wang et al., 2023c; Wu et al., 2023; Xu et al., 2024; Wang et al., 2018; Zhou et al., 2023b; Eldele et al., 2024; Duan et al., 2024)
Auxiliary Feature Extraction (§$\S$4.3) Shapelet and Motif Feature Modeling (Liang et al., 2023b; Qu et al., 2024)
Contextual Modeling and LLM Alignment (Chen et al., 2019; Anand and Nayak, 2021; Kim et al., 2023a; Choi and Kang, 2023; Rahman et al., 2021; Zhou et al., 2023a; Lin et al., 2024; Zhou et al., 2023b; Liu et al., 2019; Li et al., 2022; Lee et al., 2022b; Li et al., 2024; Luetto et al., 2023; Chowdhury et al., 2022)
Continuous Temporal and Irregular Time-Series Modeling (§$\S$4.4) Neural Differential Equations (Jhin et al., 2021; Chen et al., 2023; Abushaqra et al., 2022; Jhin et al., 2022; Rubanova et al., 2019b; Oh et al., 2024)
Implicit Neural Representations (Naour et al., 2023; Fons et al., 2022)
Auxiliary Networks (Ma et al., 2019; Zhang et al., 2023a; Bianchi et al., 2019; Schirmer et al., 2022; Shukla and Marlin, 2021; Ansari et al., 2023; Sun et al., 2021; Senane et al., 2024)
Learning-Focused Approaches (Learning Objective) Task-Adaptive Losses (§$\S$5.1) (Ma et al., 2019; Rahman et al., 2021; Hadji et al., 2021; Zhong et al., 2023; Wang et al., 2024c; Gao et al., 2024)
Non-Contrasting Losses for Temporal Dependency Learning (§$\S$5.2) Input Reconstruction Learning (Chen et al., 2019; Sanchez et al., 2019; Nguyen et al., 2018; Li et al., 2023; Yuan et al., 2019; Chen et al., 2024; Lee et al., 2024a)
Masked Prediction Learning (Zerveas et al., 2021; Biloš et al., 2023; Dong et al., 2023; Chowdhury et al., 2022; Zhang et al., 2024; Senane et al., 2024)
Customized Pretext Tasks (Zhang et al., 2022a; Guo et al., 2022; Haresh et al., 2021; Wu et al., 2018; Liang et al., 2022; Wang et al., 2020, 2021; Nguyen et al., 2018; Duan et al., 2022; Lee et al., 2022b; Fang et al., 2023; Fraikin et al., 2023; Sun et al., 2023; Dong et al., 2024; Liu et al., 2024; Bian et al., 2024)
Contrasting Losses for Consistency Modeling (§$\S$5.3) Subseries Consistency (Behrmann et al., 2021; Qian et al., 2022; Wang et al., 2020; Qian et al., 2021; Somaiya et al., 2022; Franceschi et al., 2019)
Temporal Consistency (Morgado et al., 2020; Wang et al., 2024b; Chen et al., 2022; Zhang et al., 2023b; Haresh et al., 2021; Eldele et al., 2021; Yang et al., 2022b; Tonekaboni et al., 2021; Hajimoradlou et al., 2022; Yang et al., 2023a)
Contextual Consistency (Yue et al., 2022; Zhang et al., 2023f; Zhong et al., 2023; Xu et al., 2023; Chen et al., 2022; Shin et al., 2023; Jiao et al., 2020; Chowdhury et al., 2023)
Transformation Consistency (Jenni and Jin, 2021; Eldele et al., 2021; Chen et al., 2021b)
Hierarchical and Cross-Scale Consistency (Liang et al., 2023b; Nguyen et al., 2023; Kong et al., 2020; Qing et al., 2022; Wang et al., 2023a; Zhang and Crandall, 2022; Liu et al., 2022; Dave et al., 2022; Duan et al., 2024)
Cross-Domain and Multi-Task Consistency (Zeng et al., 2021; Ding et al., 2022; Kim et al., 2023a; Han et al., 2020; Choi and Kang, 2023; Zhang et al., 2022b; Liu et al., 2023; Liu and Chen, 2023; Lee et al., 2024b; Li et al., 2024)
@@
[…]
Quantitative summary of the selected papers.
Quantitative Summary. Given the above keywords and inclusion criteria, we selected 127 papers in total. Fig. 3 shows the quantitative summary of the paper selected for review. We can notice from Fig. 3(a) that neural architectures and learning objectives are similarly considered important in designing state-of-the-art methods. Most papers were published at ICLR, NeurIPS, followed by AAAI and ICML (Fig. 3(b)). According to Fig. 3(c), we expect more papers on this topic will be published in the future.